OpenStudy (anonymous):

Imagine that for a certain person, the number of cells in a tumor doubles every three months. Supposing the tumor stars as a single cell, how many cells will there be after 3 years if this rate is maintained?

OpenStudy (anonymous):

t/3 where t is the number of total months gives the number of three month periods. (initial cells)*(2)^(t/3)=total cells. 1*2^(36/3)=2^12=4096 cells

OpenStudy (anonymous):

that looks about right thankyou